Find the centre of curvature of astroid
x^2/3+y^2/3=a^2/3
1
Expert's answer
2020-11-24T17:13:57-0500
"Solution"


Center of curvature is denoted by "P"

Lets determine the coordinates of the center of curvature at the point "P" on the astroid


Let "\theta = \frac{\pi}{6}"


"\rho = \frac{\delta s}{\delta \psi}=12\ sin\ \psi\ cos\ \psi = 6\ sin\ \psi\\
At\ \theta_p=\frac{\pi}{6},\\
\psi_p=-\frac{\pi}{6},\\
\rho_p=6\ sin\ (-\frac{\pi}{3})=6\ (-\frac{\sqrt{3}}{2})=-3 \sqrt{3}"

Now

"(X,Y)=(x_p, y_p) +\rho n"

So written

"\theta=\frac{\pi}{6},\ \psi= -\frac{\pi}{6}, x_p=\frac{3 \sqrt{3}}{2}, y_p=\frac12, \rho_p=-3 \sqrt{3}\\
n_p=(-sin(-\frac{\pi}{6})cos(-\frac{\pi}{6}))=(sin(\frac{\pi}{6})cos(\frac{\pi}{6}))= (\frac12, \frac{\sqrt{3}}{2})\\
Thus;\\
\implies(X,Y)=(\frac{3 \sqrt{3}}{2}, \frac12)+|-3 \sqrt{3}|(\frac12,\frac{\sqrt{3}}{2})\\
\implies(X,Y)=(\frac{3 \sqrt{3}}{2}, \frac12)+(-\frac{3 \sqrt{3}}{2},\frac{9}{2})\\
\therefore\ (X,Y)=(3 \sqrt {3},\ 5)"


Hence, the center of curvature of astroid is "(3 \sqrt {3},\ 5)"